I need help please. My son is needing help and it's been so long since I've done this.
TELL WHETHER THE ORDERED PAIR IS A SOLUTION OF Y=X-3; (1, -4)
I'm not sure if I write the equation as 1=-4-3 or -4=1-3 or if that is how I'm suppose to write it at all.

Answer by stanbon(75887)

TELL WHETHER THE ORDERED PAIR IS A
SOLUTION OF Y=X-3; (1, -4)
----------------------
y = x-3
-4 = 1-3
-4 = -2
No
-----
The point is not a solution of y = x-3
======================================
Cheers,
Stan H.